# UnsafeBench: Benchmarking Image Safety Classifiers on Real-World and AI-Generated Images

### v2

@@ Abstract
With the advent of text-to-image models and concerns about their misuse, developers are increasingly relying on image safety classifiers to moderate their generated unsafe images. Yet, the performance of current image safety classifiers remains unknown for both real-world and AI-generated images. In this work, we propose UnsafeBench, a benchmarking framework that evaluates the effectiveness and robustness of image safety classifiers, with a particular focus on the impact of AI-generated images on their performance. First, we curate a large dataset of 10K real-world and AI-generated images that are annotated as safe or unsafe based on a set of 11 unsafe categories of images (sexual, violent, hateful, etc.). Then, we evaluate the effectiveness and robustness of five popular image safety classifiers, as well as three classifiers that are powered by general-purpose visual language models. Our assessment indicates that existing image safety classifiers are not comprehensive and effective enough to mitigate the multifaceted problem of unsafe images. Also, there exists a distribution shift between real-world and AI-generated images in image qualities, styles, and layouts, leading to degraded effectiveness and robustness. Motivated by these findings, we build a comprehensive image moderation tool called PerspectiveVision, which addresses the main drawbacks of existing classifiers with improved effectiveness and robustness, especially on AI-generated images. UnsafeBench and PerspectiveVision can aid the research community in better understanding the landscape of image safety classification in the era of generative AI. [^1] Our webpage is at https://unsafebench.github.io/.
@@ Introduction
Parallel to real-world unsafe images, recently emerged text-to-image models like Stable Diffusion [57] and Midjourney [11] are also capable of producing unsafe images. According to a recent study [53], if intentionally misled by malicious users, text-to-image models have a 16%-51% probability of generating unsafe content like sexual, disturbing, violent content, etc. A notable real-world incident in January 2024 is the viral spread of AI-generated sexually explicit images of Taylor Swift [2], which further highlights the risks of these AI products. To mitigate these risks, AI practitioners, again, rely on image safety classifiers to identify unsafe images and block them before presenting them to end-users. For example, Stable Diffusion implements a safety filter [19] that defines 20 sensitive concepts, such as “nude,” “sex,” and “porn,” to prevent the generation of unsafe images. Additionally, extensive research [59, 35, 73, 28] applies Q16 [60] and NudeNet [15] to evaluate the safety of text-to-image models and propose new safety mechanisms.
Despite the wide use of these image safety classifiers, their performance lacks a thorough evaluation across both real-world and AI-generated unsafe images. It remains unclear if they are truly capable of detecting a broad range of unsafe content and whether their performance, i.e., effectiveness and robustness, can be generalized to AI-generated images, given that most are trained on real-world image datasets. Furthermore, the recent development of large visual language models (VLMs), such as LLaVA [45] and GPT-4V [4], provides new possibilities in this field. These models are general-purpose visual language models that have undergone rigorous safety alignment, which empowers them to understand unsafe content portrayed in images. Given the current state, it is still unknown if these advanced VLMs will supersede existing image safety classifiers with better performance.
@@
Given the lack of a comprehensive dataset covering all categories of unsafe content, our first step is curating a dataset of potentially unsafe images from public databases, including the LAION-5B [62] dataset for real-world images and the Lexica [9] website for AI-generated images. The UnsafeBench dataset is meticulously annotated and contains 10K images that encompass 11 unsafe categories and two sources (real-world or AI-generated). Next, we collect five common image safety classifiers (Q16 [60], MultiHeaded [53], SD_Filter [55], NSFW_Detector [14], and NudeNet [15]) as conventional classifiers and three VLM-based classifiers built on LLaVA [45], InstructBLIP [30], and GPT-4V [4], combined with a RoBERTa model to classify VLM’s responses. We then examine the specific unsafe content covered by these classifiers and align them with our unsafe image taxonomy. Finally, we assess these classifiers’ effectiveness with the annotated dataset and robustness against random and adversarial perturbations. We particularly focus on the varying performances of classifiers on real-world and AI-generated images. Through clustering techniques and case studies, we identify distinct characteristics in AI-generated images, such as artistic representation and grid layout, and assess whether they could interrupt the prediction from classifiers, especially those trained on real-world images.
Based on the insights drawn from UnsafeBench, we introduce PerspectiveVision, a LoRA fine-tuned LLaVA that provides fine-grained classifications of unsafe images with enhanced effectiveness and robustness. We train and evaluate PerspectiveVision on our annotated dataset and further assess its generalizability across five external datasets containing both real-world and AI-generated images.
@@
  • Real-World vs. AI-Generated Images. There exists a distribution shift between AI-generated and real-world images in image qualities, styles, and layouts. This shift leads to the degraded effectiveness for conventional classifiers trained only on real-world images, e.g., NSFW_Detector and NudeNet. AI-generated images often have unique characteristics, such as artistic representations of unsafe content and the composition of unsafe images in a grid layout, which disrupt the classifier’s predictions. Furthermore, the distribution shift also makes most classifiers more vulnerable to adversarial attacks when using AI-generated images compared to their real-world counterparts. Under the same perturbation constraint, these classifiers present lower confidence scores and higher loss increases for AI-generated images.
@@
  • PerspectiveVision. PerspectiveVision is designed to address the main drawbacks of current image safety classifiers: the most effective model (GPT-4V) being closed-source, imbalanced effectiveness, and degraded performance on AI-generated images. By including a large number of AI-generated images in the training data, PerspectiveVision achieves the highest effectiveness across five external datasets and also significantly improves robustness against various adversarial attacks for both real-world and AI-generated images.
Contributions. Our work makes three important contributions. First, we contribute to the research community with a comprehensive image dataset for image safety research. The dataset consists of 10K real-world and AI-generated images, covering unsafe content from 11 categories, with each image being meticulously annotated by three authors. It serves as a foundation dataset for future research in this field. Second, we take the first step of benchmarking the effectiveness and robustness of current image safety classifiers, particularly focusing on the impact of AI-generated images. Our assessment highlights the challenge that AI-generated unsafe images pose to existing classifiers, reducing not only their effectiveness but also their robustness, particularly against adversarial perturbations. These findings emphasize the importance for platform moderators and model developers to include AI-generated images in their training data to learn AI-specific features like artistic representation and grid layout. Finally, we contribute to the community with PerspectiveVision, which identifies a wide range of unsafe images across fine-grained categories with enhanced effectiveness and robustness on both real-world and AI-generated images. PerspectiveVision is made available as an open-source tool, establishing a baseline to detect general unsafe images. To aid future research, we are committed to responsibly sharing and continuously updating our annotated dataset and PerspectiveVision.
Ethical Considerations. The ethical concern in this study mainly lies in the manual annotation and future release of UnsafeBench images. We have undergone an ethical review by our ethical review board. Our ERB has approved the study and states that there are no ethical considerations if annotators are not exposed to images that are illegal to view or own, such as child sexual abuse materials, which do not exist in our dataset. At the same time, to mitigate any potential harm from the exposure of harmful content, we undertook all annotations ourselves, hence not exposing harmful content to other people. Regarding dataset release, we have also given careful consideration to sharing our dataset responsibly. On the one hand, sharing such a dataset will raise ethical concerns as it contains realistic unsafe images. On the other hand, dataset sharing is important to ensure reproducibility. To account for both ethical concerns and reproducibility, we will make the dataset available but only upon request and for research purposes.
@@
Classifier Backbone Training Paradigm Training Dataset
Q16 CLIP Prompt Learning SMID [29] (R)
MultiHeaded CLIP Linear Probing MultiHeaded-D [53] (A)
SD_Filter CLIP - -
NSFW_Detector CLIP Linear Probing LAION-400M Subset (R)
NudeNet Xception Fully Supervised NudeNet-D [16] (R)
Image Safety Classification Models
@@
Figure 1: Overview of UnsafeBench.
Dataset Construction
@@
Image Collection. We rely on open large-scale multimodal datasets, LAION-5B [62] and Lexica [9], to collect real-world and AI-generated unsafe images. LAION-5B is currently the largest public image-text dataset in the world [7], containing 5.85 billion image-text pairs collected from webpages [62]. We regard LAION-5B as the source to collect real-world images. Lexica [9] is an AI-generated image gallery that allows users to create and disclose images generated by models like Stable Diffusion [57]. Both datasets serve as a search engine and enable users to find the most relevant images based on a textual description. Inspired by this, we use unsafe keywords to query these datasets and collect potentially unsafe images.
@@
To collect unsafe keywords, we split the definitions of 11 unsafe categories (see Section 2) as initial unsafe keywords. Additionally, we employ a large language model, Vicuna-33b, [^3] https://huggingface.co/lmsys/vicuna-33b-v1.3. to augment the definition of each unsafe category and provide 10 specific examples. We show the generated examples in Table 8 in the Appendix. Combining the initial unsafe keywords and those generated with Vicuna-33b, we obtain 158 unsafe keywords, covering 11 unsafe categories (each containing 12-16 keywords). We then query LAION-5B and Lexica using these keywords and collect the most relevant images. After removing duplicates, we collect a total of 12,932 potentially unsafe images, comprising 5,815 images from LAION-5B and 7,117 from Lexica.
@@
Image Annotation. We perform a human annotation to determine if these collected images are truly unsafe. Three authors of this paper serve as annotators and manually annotate them independently. We require these annotators to first read the definition from each unsafe category as the criterion in determining whether an image is safe, unsafe, or not applicable (N/A). Note that N/A mainly represents noise images that cannot be identified as safe or unsafe, e.g., a blurry image or one with unidentifiable texts. The annotation process undergoes two rounds. In the first round, two annotators independently assign a label of safe, unsafe, or N/A to each image. For images where two annotators disagree, we further introduce the third annotator to provide additional labels. The final label for each image is then determined based on the majority vote among these labels. In the second round, annotators revisit images annotated as unsafe and determine the unsafe category. If an image displays a mix of unsafe elements, we choose the predominant category it violates. Through two rounds of annotation, we ensure that each image is labeled as safe, unsafe, or N/A, and if it is unsafe, it can be classified into a specific unsafe category. To evaluate the reliability of our annotation, we calculate the Fleiss’ Kappa score [33, 32], a statistical measure for assessing the agreement among annotators. The score is 0.47, indicating a moderate level of agreement among annotators [33, 32].
@@
Dataset Statistics. We show the statistics of our annotated images in Table 2 and Table 9. Overall, we annotate 12,932 images, including 6,098 safe images, 4,048 unsafe images, and 2,786 N/A images. For the following experiments, we exclude the N/A images and use the remaining images (safe + unsafe), collectively referred to as the UnsafeBench dataset. To the best of our knowledge, it is currently the most comprehensive annotated dataset in the image safety domain.
Visualizing Distribution Shift. We compare the difference between real-world and AI-generated images in image semantics and qualities, such as signal-to-noise ratio (SNR), edge density, and entropy. Figure 11 and Figure 13 show the different distribution of two groups of images. We find that a distribution shift is present in image semantics, e.g., image embeddings are separated in the 2D space between real-world and AI-generated images. The shift also exists in image qualities, e.g., AI-generated images have lower SNR and higher entropy, indicating they tend to be noisier and more complex.
@@
Given that VLMs’ outputs are sometimes lengthy sentences, they do not directly categorize into safe or unsafe classes. To convert these VLMs to classifiers, we fine-tune a language model, RoBERTa [46], to categorize LLM’s outputs into one of three classes: safe, unsafe, or uncertain. The uncertain class describes a set of outputs that do not yield a clear prediction. For example, InstructBLIP occasionally describes the content of an image without making a prediction. Note, we categorize the cases where VLMs refuse to answer as the unsafe class, as input images have triggered their internal safeguards. We collect several example outputs [^4] For example, we use “safe,” “the image is safe,” “the image is not considered unsafe as…” as example outputs for the safe class. for each class and fine-tune RoBERTa in a supervised manner. Ultimately, a VLM-based classifier consists of a VLM and the RoBERTa classifier, which function sequentially.
@@
Evaluation Methodology
Effectiveness. We evaluate the effectiveness of eight image safety classifiers (five conventional classifiers and three VLM-based classifiers). For conventional classifiers, we feed images from each unsafe category and obtain binary predictions (safe/unsafe). Regarding VLM-based classifiers, we input both images with the designed prompts to gather outputs, which are then classified into safe/unsafe/uncertain classes by the fine-tuned RoBERTa. Uncertain predictions account for about 0.1%. However, since conventional classifiers directly predict images as safe or unsafe, to ensure a fair comparison, we randomly assign a prediction of safe or unsafe to VLM’s uncertain predictions. Recall that we design three prompts for VLMs, resulting in three separate predictions for each image. The final prediction is determined by a majority vote.
@@
Robustness. We evaluate the robustness of classifiers using images perturbed with both random noise (Gaussian noise) and adversarial noise (i.e., adversarial examples) [49, 47]. Adversarial examples ($x_{adv}$) are original inputs ($x$) with optimized perturbations ($\Delta x$), which maximizes the loss ($L(\theta,x+\Delta x,y)$) of a model ($\theta$) and fools the model into making incorrect predictions. The p-norm of the perturbation vector is limited by a small $\epsilon$ to ensure the perturbation is imperceptible.
However, for VLMs, directly solving Equation 1 does not necessarily create adversarial examples that lead to opposite predictions. Unlike binary classifiers, drifting the VLM away from its original predictions, e.g., “the image is safe,” could result in unexpected outputs that are unrelated to the image safety classification task. To address this, we transform the untargeted attack in Equation 1 into targeted attacks (targeting the opposite class), which are equivalent in binary classification settings. For example, if the VLM initially classifies an image as safe, we then optimize the perturbation such that the output moves toward the “unsafe” direction by setting the target output as “unsafe.” Therefore, instead of maximizing the loss between the classifier’s prediction and the original label ($y$), here, we minimize the loss of a VLM between its prediction and the defined target output ($y_{tar}$), as shown in Equation 2. By solving the equation, we update optimized perturbations until the RoBERTa classifier classifies the VLM output as the opposite class from the image label. Using this strategy, we create adversarial examples using FGSM, PGD, and DeepFool on open-source VLM-based classifiers (LLaVA and InstructBLIP). Note that we exclude GPT-4V from this evaluation because crafting adversarial examples requires model gradients, which is not available for GPT-4V.
To maintain the same perturbation budget for both conventional classifiers and VLMs, we use the $L_{\infty}$ norm, set $\epsilon$ to 0.01, and also limit the number of optimization iterations to a maximum of 100. We demonstrate the perturbed images and different types of perturbations in Figure 2.
Evaluation Results
Dataset Classifier Hate Harassment Violence Self-Harm Sexual Shocking Illegal Activity Deception Political Health Spam Overall
LAION-5B (Real-World) Q16 0.641 0.517 0.693 0.421 - 0.630 0.681 0.762 0.271 0.144 - -
MultiHeaded 0.320 - 0.247 0.000 0.692 0.644 - - 0.209 - - -
SD_Filter - - - - 0.833 - - - - - - -
NSFW_Detector - 0.517 - - 0.783 - - - - - - -
NudeNet - - - - 0.650 - - - - - - -
LLaVA* 0.227 0.167 0.413 0.451 0.674 0.714 0.498 0.376 0.116 0.333 0.059 0.426
InstructBLIP* 0.351 0.394 0.606 0.275 0.796 0.467 0.722 0.653 0.444 0.379 0.380 0.520
GPT-4V* 0.556 0.706 0.774 0.557 0.866 0.724 0.897 0.827 0.605 0.405 0.718 0.712
Lexica (AI-Generated) Q16 0.336 0.416 0.612 0.521 - 0.836 0.497 0.384 0.597 0.615 - -
MultiHeaded 0.225 - 0.533 - 0.815 0.780 - - 0.744 - - -
SD_Filter - - - - 0.727 - - - - - - -
NSFW_Detector - 0.259 - - 0.678 - - - - - - -
NudeNet - - - - 0.596 - - - - - - -
LLaVA* 0.169 0.224 0.632 0.663 0.866 0.826 0.534 0.179 0.340 0.665 0.045 0.634
InstructBLIP* 0.178 0.332 0.629 0.383 0.757 0.795 0.665 0.302 0.785 0.663 0.519 0.589
GPT-4V* 0.254 0.635 0.712 0.613 0.827 0.875 0.701 0.422 0.909 0.621 0.171 0.707
Overall Q16 0.533 0.475 0.648 0.483 - 0.784 0.571 0.652 0.482 0.498 - -
MultiHeaded 0.292 - 0.426 - 0.757 0.749 - - 0.600 - - -
SD_Filter - - - - 0.785 - - - - - - -
NSFW_Detector - 0.449 - - 0.738 - - - - - - -
NudeNet - - - - 0.624 - - - - - - -
LLaVA* 0.210 0.189 0.550 0.590 0.780 0.800 0.519 0.332 0.266 0.575 0.054 0.549
InstructBLIP* 0.270 0.368 0.615 0.333 0.777 0.697 0.687 0.506 0.660 0.545 0.490 0.559
GPT-4V* 0.423 0.681 0.738 0.590 0.847 0.839 0.780 0.673 0.780 0.492 0.537 0.709
Effectiveness
@@
Conventional vs. VLM-Based Classifiers. The top-performing image safety classifier is the commercial VLM-based classifier, GPT-4V. It achieves the highest overall F1-Score, surpassing LLaVA and InstructBLIP, as well as other conventional classifiers in most unsafe categories, except for Hate and Deception. GPT-4V shows exceptional effectiveness in detecting Sexual (0.847), Shocking (0.839), Political (0.780), Illegal Activity (0.780), and Violence content (0.738). Among conventional classifiers, Q16 stands out by identifying the broadest spectrum of unsafe content, with an F1-Score from 0.475 to 0.784. However, Q16 does not support the detection of Sexual content and requires improvements in Harassment, Self-Harm, and Political categories. In contrast, MultiHeaded, SD_Filter, NSFW_Detector, and NudeNet can detect sexual images with an F1-Score of 0.624-0.785.

### v3

@@ Abstract
With the advent of text-to-image models and concerns about their misuse, developers are increasingly relying on image safety classifiers to moderate their generated unsafe images. Yet, the performance of current image safety classifiers remains unknown for both real-world and AI-generated images. In this work, we propose UnsafeBench, a benchmarking framework that evaluates the effectiveness and robustness of image safety classifiers, with a particular focus on the impact of AI-generated images on their performance. First, we curate a large dataset of 10K real-world and AI-generated images that are annotated as safe or unsafe based on a set of 11 unsafe categories of images (sexual, violent, hateful, etc.). Then, we evaluate the effectiveness and robustness of five popular image safety classifiers, as well as three classifiers that are powered by general-purpose visual language models. Our assessment indicates that existing image safety classifiers are not comprehensive and effective enough to mitigate the multifaceted problem of unsafe images. Also, there exists a distribution shift between real-world and AI-generated images in image qualities, styles, and layouts, leading to degraded effectiveness and robustness. Motivated by these findings, we build a comprehensive image moderation tool called PerspectiveVision, which improves the effectiveness and robustness of existing classifiers, especially on AI-generated images. UnsafeBench and PerspectiveVision can aid the research community in better understanding the landscape of image safety classification in the era of generative AI. [^1] The project’s webpage is https://unsafebench.github.io/.
@@ Introduction
Parallel to real-world unsafe images, AI-generated unsafe images are becoming a new threat [53, 2]. According to a recent study [53], AI models like Stable Diffusion [57] could have a 16%-51% probability of generating unsafe content like sexual, disturbing, violent content, etc, when intentionally misled. To mitigate these risks, AI practitioners, again, rely on image safety classifiers to identify unsafe images and block them before presenting them to end-users. For example, Stable Diffusion implements a safety filter [18] that defines 20 sensitive concepts, such as “nude,” “sex,” and “porn,” to prevent the generation of unsafe images. Additionally, extensive research [59, 34, 70, 27] applies Q16 [60] and NudeNet [14] to evaluate the safety of text-to-image models and propose new safety mechanisms.
Despite the wide use of these image safety classifiers, their performance lacks a thorough evaluation across both real-world and AI-generated unsafe images. Particularly for AI-generated unsafe images, due to differences in visual styles and distributions, it is unknown if current image safety classifiers can still effectively identify these AI products, given that most are trained on real-world image datasets. Additionally, as supervised classifiers are vulnerable to adversarial attacks, it is unclear whether they can maintain robustness across both real-world and AI-generated unsafe content. Furthermore, the recent development of large visual language models (VLMs), such as LLaVA [45] and GPT-4V [5], provides new possibilities in this field. These models are general-purpose visual language models that have undergone rigorous safety alignment, which empowers them to understand unsafe content portrayed in images. Given the current state, it is still unknown if these advanced VLMs will supersede existing image safety classifiers with better performance.
Our Work. To address these concerns, we conduct a thorough evaluation to understand the performance of current image safety classifiers and VLMs in identifying unsafe content. Our evaluation examines three critical aspects: (1) the performance of conventional classifiers vs. VLMs, (2) the performance on real-world vs. AI-generated images, and (3) the performance of effectiveness across various unsafe categories and robustness under adversarial attacks.
@@
Given the lack of a comprehensive dataset covering all categories of unsafe content, our first step in the UnsafeBench framework is curating a dataset of potentially unsafe images from public databases, including the LAION-5B [61] dataset for real-world images and the Lexica [9] website for AI-generated images. The UnsafeBench dataset is meticulously annotated and contains 10K images that encompass 11 unsafe categories and two sources (real-world or AI-generated). Next, we collect five common image safety classifiers (Q16 [60], MultiHeaded [53], SD_Filter [55], NSFW_Detector [13], and NudeNet [14]) as conventional classifiers and three VLM-based classifiers built on LLaVA [45], InstructBLIP [29], and GPT-4V [5], combined with a RoBERTa model to classify VLM’s responses. We then examine the specific unsafe content covered by these classifiers and align them with our unsafe image taxonomy. Finally, we assess these classifiers’ effectiveness with the annotated dataset and robustness against random and adversarial perturbations. We particularly focus on the varying performances of classifiers on real-world and AI-generated images. Through clustering techniques and case studies, we identify specific characteristics in AI-generated images, such as artistic representation and grid layout, and assess whether they could interrupt the prediction from classifiers, especially those trained on real-world images.
Additionally, facing the AI threats, we take the first step at building an image moderation tool that could generalize to AI-generated unsafe images with enhanced effectiveness and robustness, PerspectiveVision. It is a LoRA fine-tuned LLaVA that provides fine-grained classifications of unsafe images under the user-defined unsafe taxonomy. We train and evaluate PerspectiveVision on our annotated dataset and further assess its generalizability across multiple external datasets containing both real-world and AI-generated images.
@@
  • Real-World vs. AI-Generated Images. There exists a distribution shift between AI-generated and real-world images in image qualities, styles, and layouts. This shift leads to potential degraded effectiveness for conventional classifiers trained only on real-world images, e.g., NSFW_Detector and NudeNet. AI-generated images often have specific characteristics, such as artistic representations of unsafe content and grid layout, which might disrupt the classifier’s predictions. Furthermore, the distribution shift also makes most classifiers more vulnerable to adversarial attacks when using AI-generated images compared to their real-world counterparts. Under the same perturbation constraint, these classifiers present lower confidence scores and higher loss increases for AI-generated images.
@@
  • PerspectiveVision. PerspectiveVision is designed to improve the effectiveness and robustness of current image safety classifiers on AI-generated images. By training LLaVA on a large number of AI-generated images, PerspectiveVision achieves the highest overall F1-Score across six evaluation datasets, including one in-distribution and five external out-of-distribution datasets. It also significantly improves robustness against various adversarial attacks for both real-world and AI-generated images.
Contributions. Our work makes three important contributions:
  1. 1.
    First, we contribute to the research community with a comprehensive image dataset for image safety research. The dataset consists of 10K real-world and AI-generated images, covering unsafe content from 11 categories, with each image being meticulously annotated by three authors. It serves as a valuable foundation dataset for future research relevant to AI-generated unsafe content.
  2. 2.
    Second, we take the first step of benchmarking the effectiveness and robustness of current image safety classifiers, particularly focusing on the impact of AI-generated images. Our assessment highlights the challenge that AI-generated unsafe images pose to existing classifiers, potentially reducing not only their effectiveness but also their robustness, particularly against adversarial perturbations. Since AI-generated unsafe images consistently make most classifiers, including a VLM (LLaVA) more vulnerable to adversarial attacks, there is a possibility for higher successful jailbreak attacks against VLMs, using AI-generated unsafe visuals. These findings emphasize the importance for platform moderators and model developers to include AI-generated images in their training data to learn AI-specific features and improve both effectiveness and robustness.
  3. 3.
    Finally, we contribute to the community with PerspectiveVision, which identifies a wide range of unsafe images across fine-grained categories with enhanced effectiveness and robustness on both real-world and AI-generated images. PerspectiveVision is made available as an open-source tool, establishing a baseline to detect (AI-generated) unsafe images.
Ethical Considerations. We have undergone an ethical review by our institution’s Ethics Review Board (ERB). Our ERB has approved the study and states that there are no ethical considerations if annotators are not exposed to images that are illegal to view or own, such as child sexual abuse materials, which do not exist in our dataset.
Nonetheless, we recognize that ethical responsibility extends beyond the ERB approval. The main ethical concerns in this study involve the annotation process, demonstration of unsafe examples, and future release of UnsafeBench images. First, to minimize potential harm from exposure to harmful content, all annotations are conducted by our research team, ensuring that no external annotators are subjected to distressing material. Second, we implement strict measures, including exposure limits, scheduled breaks, and regular mental health check-ins, to ensure the annotators’ well-being. Regarding the demonstration of unsafe images, since this study involves unsafe content, displaying unsafe examples is unavoidable. However, we censor Not-Safe-For-Work (NSFW) images and avoid displaying unsafe images that might be offensive to different communities. Finally, regarding dataset release, we carefully balance ethical concerns with the need for reproducibility and will publicly release the dataset upon request for research purposes.
@@
Figure 1: High-level overview of UnsafeBench.
UnsafeBench Dataset
@@
Image Collection. We rely on open large-scale multimodal datasets, LAION-5B [61] and Lexica [9], to collect real-world and AI-generated unsafe images. LAION-5B is currently the largest public image-text dataset in the world [7], containing 5.85 billion image-text pairs collected from webpages [61]. We regard LAION-5B as the source to collect real-world images. Lexica [9] is one of the largest AI image galleries, containing over 5 million [9] AI images generated by real-world users using models like Stable Diffusion [57]. Both datasets serve as a search engine and enable users to find the most relevant images based on a textual description. Inspired by this, we use unsafe keywords to query these datasets and collect potentially unsafe images.
To collect unsafe keywords, we split the definitions of 11 unsafe categories (see Appendix Table 9) as initial unsafe keywords (1-5). For instance, the initial unsafe keywords for the Hate category are “hateful symbols,” “negative stereotypes,” “comparing certain groups to animals/objects,” and “promoting hate based on identity.” Given the limited number and the broad scope of initial keywords, we additionally employ an LLM to augment the definition of each unsafe category by providing 10 more specific examples, e.g., “Swastika,” “Confederate flag,” “anti-Semitic symbols,” for the Hate category. Specifically, we use Vicuna-33b, [^3] https://huggingface.co/lmsys/vicuna-33b-v1.3. as it is more compliant with sensitive requests like generating unsafe keywords. We show the generated examples in Table 9 in the Appendix. These keywords are manually verified by three annotators independently before the annotation process to ensure that (1) each correctly reflects the associated unsafe category, and (2) minimal overlap within and across unsafe categories. We also cross-check these augmented examples with those in the original definition. In cases of uncertain keywords, we exclude them from the unsafe keyword list.
@@
Combining the initial unsafe keywords and those augmented ones, we obtain 158 distinct unsafe keywords, covering 11 unsafe categories (each containing 12-16 keywords). We then query LAION-5B and Lexica using these keywords and collect the most relevant images. After removing duplicates, we collect a total of 12,932 potentially unsafe images, comprising 5,815 images from LAION-5B and 7,117 from Lexica.
@@
Image Annotation. We perform a human annotation to determine if these collected images are truly unsafe. Three authors of this paper serve as annotators and manually annotate them independently. We require these annotators to first read the definition from each unsafe category as the criterion in determining whether an image is safe, unsafe, or not applicable (N/A). Note that N/A mainly represents noise images that cannot be identified as safe or unsafe, e.g., a blurry image or one with unidentifiable texts. The annotation process undergoes two rounds. In the first round, two annotators independently assign a label of safe, unsafe, or N/A to each image. For images where two annotators disagree, we further introduce the third annotator to provide additional labels. The final label for each image is then determined based on the majority vote among these labels. In the second round, annotators revisit images annotated as unsafe and determine the unsafe category. If an image displays a mix of unsafe elements, we choose the predominant category it violates. Through two rounds of annotation, we ensure that each image is labeled as safe, unsafe, or N/A, and if it is unsafe, it can be classified into a specific unsafe category. To evaluate the reliability of our annotation, we calculate the agreement percentage and the Fleiss’ Kappa score [32, 31] to reflect the agreement among annotators. The agreement percentage is 0.780, indicating that the majority of annotations are consistent. Fleiss’ Kappa is 0.697, also suggesting a level of substantial agreement (0.61-0.80) among annotators [32, 31].
@@
Dataset Statistics. We show the statistics of our annotated images in Table 1 and Table 10. Overall, we annotate 12,932 images, including 6,098 safe, 4,048 unsafe, and 2,786 N/A images. For the following experiments, we exclude the N/A images and use the remaining images (safe + unsafe), collectively referred to as the UnsafeBench dataset. To the best of our knowledge, it is currently the most comprehensive annotated dataset in the image safety domain.
Visualizing Distribution Shift. We compare the difference between real-world and AI-generated images in image semantics and qualities, such as signal-to-noise ratio (SNR), edge density, and entropy. Figure 11 and Figure 12 (Appendix) show the different distributions of two groups of images. We find that a distribution shift is present in image semantics, e.g., image embeddings are separated in the 2D space between real-world and AI-generated images. The shift also exists in image qualities, e.g., AI-generated images have lower SNR and higher entropy, indicating they tend to be noisier and more complex.
@@
Given that VLMs’ outputs are sometimes lengthy sentences, they do not directly categorize into safe or unsafe classes. To convert these VLMs to classifiers, we fine-tune a language model, RoBERTa [46], to categorize LLM’s outputs into one of three classes: safe, unsafe, or uncertain. The uncertain class describes a set of outputs that do not yield a clear prediction. For example, InstructBLIP occasionally describes the content of an image without making a prediction. Note, we categorize the cases where VLMs refuse to answer as the unsafe class, as input images have triggered their internal safeguards. We provide specific training details and the reliability result of the RoBERTa classifier in Section A.2 in the Appendix. It achieves an accuracy of 0.992 and an F1-Score of 0.991 on 600 randomly sampled responses generated by three VLMs. Ultimately, a VLM-based classifier consists of a VLM and the RoBERTa classifier, which function sequentially.
@@
Dataset Classifier Hate Harassment Violence Self-Harm Sexual Shocking Illegal Activity Deception Political Health Spam
LAION-5B (Real-World) Q16 0.641 0.517 0.693 0.421 - 0.630 0.681 0.762 0.271 0.144 -
MultiHeaded 0.320 - 0.247 0.000 0.692 0.644 - - 0.209 - -
SD_Filter - - - - 0.833 - - - - - -
NSFW_Detector - 0.517 - - 0.783 - - - - - -
NudeNet - - - - 0.650 - - - - - -
LLaVA* 0.227 0.167 0.413 0.451 0.674 0.714 0.498 0.376 0.116 0.333 0.059
InstructBLIP* 0.351 0.394 0.606 0.275 0.796 0.467 0.722 0.653 0.444 0.379 0.380
GPT-4V* 0.556 0.706 0.774 0.557 0.866 0.724 0.897 0.827 0.605 0.405 0.718
Lexica (AI-Generated) Q16 0.336 0.416 0.612 0.521 - 0.836 0.497 0.384 0.597 0.615 -
MultiHeaded 0.225 - 0.533 - 0.815 0.780 - - 0.744 - -
SD_Filter - - - - 0.727 - - - - - -
NSFW_Detector - 0.259 - - 0.678 - - - - - -
NudeNet - - - - 0.596 - - - - - -
LLaVA* 0.169 0.224 0.632 0.663 0.866 0.826 0.534 0.179 0.340 0.665 0.045
InstructBLIP* 0.178 0.332 0.629 0.383 0.757 0.795 0.665 0.302 0.785 0.663 0.519
GPT-4V* 0.254 0.635 0.712 0.613 0.827 0.875 0.701 0.422 0.909 0.621 0.171
Overall Q16 0.533 0.475 0.648 0.483 - 0.784 0.571 0.652 0.482 0.498 -
MultiHeaded 0.292 - 0.426 - 0.757 0.749 - - 0.600 - -
SD_Filter - - - - 0.785 - - - - - -
NSFW_Detector - 0.449 - - 0.738 - - - - - -
NudeNet - - - - 0.624 - - - - - -
LLaVA* 0.210 0.189 0.550 0.590 0.780 0.800 0.519 0.332 0.266 0.575 0.054
InstructBLIP* 0.270 0.368 0.615 0.333 0.777 0.697 0.687 0.506 0.660 0.545 0.490
GPT-4V* 0.423 0.681 0.738 0.590 0.847 0.839 0.780 0.673 0.780 0.492 0.537
Effectiveness Assessment
Methodology
We evaluate the effectiveness of eight image safety classifiers (five conventional classifiers and three VLM-based classifiers). For conventional classifiers, we feed images from each unsafe category and obtain binary predictions (safe/unsafe). Regarding VLM-based classifiers, we input both images with the designed prompts to gather outputs, which are then classified into safe/unsafe/uncertain classes by the fine-tuned RoBERTa. Uncertain predictions account for about 0.1%. However, since conventional classifiers directly predict images as safe or unsafe, to ensure a fair comparison, we randomly assign a prediction of safe or unsafe to VLM’s uncertain predictions. Recall that we design three prompts for VLMs, resulting in three separate predictions for each image. The final prediction is determined by a majority vote.
@@
Effectiveness Result
@@
Conventional vs. VLM-Based Classifiers. The top-performing image safety classifier is the commercial VLM-based classifier, GPT-4V. It achieves the highest F1-Score in most unsafe categories, except for Hate and Deception. GPT-4V shows exceptional effectiveness in detecting Sexual (0.847), Shocking (0.839), Political (0.780), Illegal Activity (0.780), and Violence content (0.738). Among conventional classifiers, Q16 stands out by identifying the broadest spectrum of unsafe content, with an F1-Score from 0.475 to 0.784. However, Q16 does not support the detection of Sexual content and requires improvements in Harassment, Self-Harm, and Political categories. In contrast, MultiHeaded, SD_Filter, NSFW_Detector, and NudeNet can detect sexual images with an F1-Score of 0.624-0.785.